The Dahlia Medical Center has 32 labor rooms, 20 combination labor and delivery rooms, and 2 delivery rooms. All of these facilities operate around the clock. Time spent in labor rooms varies from hours to days, with an average of about a day. The average uncomplicated delivery requires about one hour in the delivery room. The average time in a combination labor-delivery room is about 24 hours.

During an exceptionally busy three-day period, 95 healthy babies were born at Dahlia Medical Center. 65 babies were born in separate labor and delivery rooms and 30 were born in combined labor and delivery rooms. Which of the facilities (labor rooms, combination labor and delivery rooms, or delivery rooms) had the greatest utilization rate?

The ______(labor rooms/ combination labor and delivery rooms/ delivery rooms)had the highest utilization rate of _______ %. (Enter your response as a percent rounded to two decimal places.)
Answer #1

Labor room capacity = Labor rooms*no of days*No of hours = 32 rooms*3 days* 24 hours = 2304

Labor room utilization = (65 babies*24hours)/2304 =0.7444 = 67.71%

Delivery room capacity = 2 rooms*24 hours*3 days= 144 hours

Delivery room utilization = 65 babies * 1 hour /144 = 45.14%

Combination labor-delivery room capacity = 20*24*3days = 1440

Combination labor-delivery room utilization = (30*24)/1440 = 50%

Labor room utilization had the highest utilization rate of 67.71%